# Generalize across Homophily and Heterophily: Hybrid Spectral Graph Pre-Training and Prompt Tuning

### v2

@@ Abstract
Graph “pre-training and prompt-tuning” aligns downstream tasks with pre-trained objectives to enable efficient knowledge transfer under limited supervision. However, existing methods rely on homophily-based low-frequency knowledge, failing to handle diverse spectral distributions in real-world graphs with varying homophily. Our theoretical analysis reveals a spectral specificity principle: optimal knowledge transfer requires alignment between pre-trained spectral filters and the intrinsic spectrum of downstream graphs. Under limited supervision, large spectral gaps between pre-training and downstream tasks impede effective adaptation. To bridge this gap, we propose the HS-GPPT model, a novel framework that ensures spectral alignment throughout both pre-training and prompt-tuning. We utilize a hybrid spectral filter backbone and local-global contrastive learning to acquire abundant spectral knowledge. Then we design prompt graphs to align the spectral distribution with pretexts, facilitating spectral knowledge transfer across homophily and heterophily. Extensive experiments validate the effectiveness under both transductive and inductive learning settings. Our code is available at https://anonymous.4open.science/r/HS-GPPT-62D2/.
@@ 1 Introduction
Graph “pre‑training and prompt tuning” (Sun et al. 2022; Liu et al. 2023; Fang et al. 2024; Sun et al. 2023; Yu et al. 2024) has recently emerged as a powerful paradigm for handling data scarcity in graph‑based learning. By freezing a pre‑trained Graph Neural Network (GNN) and injecting small, learnable prompt structures (e.g., virtual nodes, subgraphs, or modified edges) into the input, one can adapt the model pretrained on graphs using self-supervision to new tasks on a target graph without updating its full parameter set. Under limited supervision, these prompts effectively bridge the gap between downstream objectives and the original pre‑training task, boosting knowledge transfer. Despite initial success, existing prompt tuning methods rely on the homophily assumption, learning node similarities in pre-training to empower downstream tasks. However, real-world graphs often contain heterophilic edges, meaning connections between dissimilar nodes, which leads to low homophily. Recent work (Yu et al. 2024) shows that homophily-based prompt tuning methods struggle to generalize across graphs with varying homophily levels, under both transductive and inductive settings.
@@
To address this issue, we leverage the connection between homophily and spectral characteristics (Chen, Lei, and Wei 2024; Wan et al. 2024; Duan et al. 2024), and examine limitations of existing graph prompt tuning methods from a spectral perspective. As shown in Figure 1, we observe significant spectral distribution diversity across graphs with different homophily levels. Notably, lower homophily correlates with an increase in the high-frequency area $S_{high}$. However, current methods are restricted to a single kind of spectral knowledge (i.e., low-frequency information) with low-frequency filters (e.g., GCN (Kipf and Welling 2016)), and some (Sun et al. 2022; Liu et al. 2023) even utilize self-supervised tasks which only preserve low-frequency information (Liu et al. 2022). Given this observation, we further theoretically prove the spectral specificity that optimal knowledge transfer requires alignment between pre-trained spectral filters and the intrinsic spectrum of the downstream graph. While prompt tuning facilitates knowledge transfer through input manipulation, large spectral gaps between pre-trained knowledge and downstream graphs hinder optimal parameter learning due to limited supervision, leading to the failure of existing methods.
@@
To address the heterophily issue in graph prompt tuning, we propose the Hybrid Spectral Graph Pre-training and Prompt Tuning model (HS-GPPT), which overcomes two challenges. The first challenge is: What spectral knowledge should be learned, and how? Our analysis on spectral specificity shows that acquiring abundant spectral knowledge is essential for downstream generalization. Therefore, we adopt a hybrid GNN backbone based on Beta wavelet transformations (Tang et al. 2022; Xu et al. 2024), where each filter captures a distinct spectral band. Building on prior methods (Chen, Lei, and Wei 2024) that handle only low- and high-pass filters, we extend the pre-training process to cover multiple frequency bands, allowing the model to learn filter-specific knowledge and their integration.
@@
The second challenge is: How can prompts adaptively align with this abundant spectral knowledge to enhance performance and prevent negative transfer? We introduce spectral-aligned prompt graphs: lightweight, learnable subgraphs that manipulate the spectral distribution of downstream graphs to match the characteristics of the pre-trained filters. We further theoretically prove the spectral alignment principle: for each filter, an appropriately learned prompt graph can align spectral distributions, thus facilitating more effective knowledge transfer. By assigning a dedicated prompt graph to each filter, our method achieves fine-grained alignment with the pre-trained spectral knowledge.
@@
In summary, our contributions are: (i) The first theoretical analysis of hybrid spectral graph prompt tuning, proving the need for rich spectral pre-training and establishing spectral alignment principles; (ii) HS-GPPT, a novel framework leveraging abundant spectral knowledge for fine-grained downstream alignment, boosting knowledge transfer; (iii) Experiments across datasets with varying homophily show HS-GPPT outperforms baselines in both transductive and inductive settings, validating its generalization.
@@ 2 Related Work
Graph Prompt Tuning. Graph prompt tuning (Sun et al. 2022; Fang et al. 2024; Sun et al. 2023; Liu et al. 2023; Yu et al. 2024) tackles data scarcity by leveraging frozen pre-trained knowledge through carefully designed prompts. For example, GPrompt (Liu et al. 2023) uses prompt vectors to unify pre-training and downstream tasks under a common template, while GPF (Fang et al. 2024) inserts prompt nodes adaptable to various pre-training strategies. These methods, tailored to homophilic graphs, rely mainly on low-frequency signals and can be seen as special cases of our approach when restricted to low-frequency information. Our spectral analysis shows that under sparse supervision, large spectral gaps impede optimal parameter learning, leading to the failure of existing homophily-based methods. ProNoG (Yu et al. 2024) takes an initial step toward heterophilic prompt tuning but lacks a principled spectral analysis. In contrast, our method enriches pre-trained knowledge, narrows spectral gaps, and enables effective alignment and robust performance even with limited labels.
Heterophilic Graph Learning. Recent methods (He et al. 2023; Xiao et al. 2022, 2024; Bo et al. 2021; Chen, Lei, and Wei 2024; Wan et al. 2024) design GNN architectures and pre-training objectives for heterophilic graphs, such as leveraging high-frequency information (Bo et al. 2021; Chen, Lei, and Wei 2024; Wan et al. 2024) and discovering potential neighbors (Jin et al. 2021; Pei et al. 2020). However, they rely on full-model fine-tuning. Without a mechanism to align pre-trained knowledge with downstream tasks, directly using their heterophilic priors can cause negative transfer, especially under limited supervision. Thus, these models and pre-training schemes for heterophilic graphs are ill-suited for prompt tuning methods.
@@ 3.1 Preliminaries
Recognizing that node classification is the main challenge on heterophilic graphs, we follow prior work (Bo et al. 2021; Chen, Lei, and Wei 2024) and select this task as our downstream objective. Specifically, the graph prompt tuning occurs under limited supervision, where for each class, only $K$ nodes are provided, also known as $K$-shot learning.
Homophily. Homophily refers to the tendency of nodes to connect with other nodes that are similar to themselves. Edge homophily level $h$ (Zhu et al. 2020) measures the fraction of homophilic edges that connect nodes share the same label, i.e., $h=\frac{|\{e_{uv}\in{\mathcal{E}}:y_{u}=y_{v}\}|}{|\mathcal{E}|}\in[0,1]$, where $|\mathcal{E}|$ denotes the number of edges and $y_{i}$ is the label of node $v_{i}$. A larger $h$ means a larger degree of homophily.
Spectral Graph Filters. Given a graph $\mathcal{G}={\{\mathcal{V},\mathbf{X},{\mathcal{E}}}\}$, let $\mathbf{A}\in\mathbb{R}^{N\times N}$ be the adjacency matrix and $\mathbf{D}\in\mathbb{R}^{N\times N}$ denote the diagonal degree matrix with $\mathbf{D}_{ii}=\sum_{j}\mathbf{A}_{ij}$. The normalized graph Laplacian matrix is $\mathbf{L}=\mathbf{I}-\mathbf{D}^{-\frac{1}{2}}\mathbf{A}\mathbf{D}^{-\frac{1}{2}}$, where $\mathbf{I}$ is an identity matrix. $\mathbf{L}$ can be eigendecomposed as $\mathbf{L}=\mathbf{U}\mathbf{\Lambda}\mathbf{U}^{T}$, where $\mathbf{U}=[\mathbf{u}_{1},\mathbf{u}_{2},...,\mathbf{u}_{N}]$ are orthonormal eigenvectors and $\mathbf{\Lambda}=diag([\lambda_{1},\lambda_{2},...,\lambda_{N})]$ are corresponding eigenvalues. Typically, $\mathbf{U}$ is the graph Fourier basis and $\mathbf{\Lambda}$ are the frequencies. Given $\lambda_{i}\ \in[0,2]$, $\lambda_{i}$ near 2 are high frequencies, while $\lambda_{i}$ near 0 are low frequencies. The objective of spectral graph filters is to design a function $g(\cdot)$ on $\mathbf{\Lambda}$ to learn the graph representation $\mathbf{Z}$:
@@
From a spatial view, different filtering characteristics aggregate diverse neighbor information. Low-pass filters capture node-neighbor similarity, while high-pass filters capture differences (Luo et al. 2024).
Suppose the feature dimension $d=1$, the nodes’ feature is denoted as $\mathbf{x}\in\mathbb{R}^{N\times 1}$. The spectral graph signal $\mathbf{\hat{x}}$ is obtained by $\mathbf{\hat{x}}=(\hat{x}_{1},\hat{x}_{2},...,\hat{x}_{N})^{T}=\mathbf{U}^{T}\mathbf{x}$, where $\hat{x}_{i}=\mathbf{u}_{i}^{T}\mathbf{x}$ gives the projection of $\mathbf{x}$ to frequency $\lambda_{i}$. We denote the spectral energy as $\hat{x}_{k}^{2}/\sum_{i=1}^{N}\hat{x}_{i}^{2}$, which reflects the spectral distribution. The larger the $\hat{x}_{k}^{2}/\sum_{i=1}^{N}\hat{x}_{i}^{2}$ is, the more frequency components corresponding to $\lambda_{i}$ will be. To observe the spectral distribution intuitively, we introduce the high-frequency area  (Tang et al. 2022).
@@
For a graph with spectral graph signal $\mathbf{\hat{x}}=(\hat{x}_{1},\hat{x}_{2},...,\hat{x}_{N})^{T}=\mathbf{U}^{T}\mathbf{x}$, the high-frequency area is calculated as $S_{high}=\frac{\sum_{k=1}^{N}\lambda_{k}\hat{x}_{k}^{2}}{\sum_{k=1}^{N}\hat{x}_{k}^{2}}=\frac{\mathbf{x}^{T}\mathbf{L}\mathbf{x}}{\mathbf{x}^{T}\mathbf{x}}$. $S_{high}$ quantifies the spectral distribution ($S_{high}\in[0,2]$). A larger $S_{high}$ indicates stronger high-frequency components, while a smaller value implies greater low-frequency components.
@@ 3.2 Theoretical Analysis
In this section, we first theoretically analyze and experimentally verify the spectral distribution diversity of real-world graphs. We then prove the spectral specificity principle: optimal knowledge transfer requires alignment between pre-trained spectral filters and the downstream graph’s intrinsic spectrum. Thus, a single graph filter is hard to align with the spectral distribution diversity, especially under limited supervision, highlighting the need for abundant spectral pre-trained knowledge.
Spectral Distribution Diversity.
Given a graph $\mathcal{G}$ with node label $\mathbf{Y}$ and homophily level $h$, we denote the expectation of normalized intra-class and inter-class distances (Ju et al. 2025) as:
$\mathbb{E}(d_{intra})=\mathbb{E}\left[\frac{(x_{u}-x_{v})^{2}}{\mathbf{x}^{T}\mathbf{x}}\right]_{e_{uv}\in\mathcal{E},y_{u}=y_{v}}$ (3)
$\mathbb{E}(d_{inter})=\mathbb{E}\left[\frac{(x_{u}-x_{v})^{2}}{\mathbf{x}^{T}\mathbf{x}}\right]_{e_{uv}\in\mathcal{E},y_{u}\neq y_{v}}$ (4)
$\mathbb{E}(d_{intra})$ and $\mathbb{E}(d_{inter})$ reflect the node feature distribution in the graph. In general, the distance between nodes of different classes is larger than that between nodes of the same class. Hence, $\mathbb{E}(d_{inter})>\mathbb{E}(d_{intra})$. With the above definition, we propose Theorem 3.2 to formally depict the correlation between spectral distribution and homophily.
Theorem 3.2.
Given a graph $\mathcal{G}$ with homophily level $h$. We denote the normalized intra-class and inter-class distances as $\mathbb{E}(d_{intra})$ and $\mathbb{E}(d_{inter})$, respectively. Then
$S_{high}=\frac{{|\mathcal{E}|}}{2}\cdot\Big{\{}\mathbb{E}(d_{inter})-\big{[}\mathbb{E}(d_{inter})-\mathbb{E}(d_{intra})\big{]}\cdot h\Big{\}}.$ (5)
Generally, we have $\mathbb{E}(d_{inter})>\mathbb{E}(d_{intra})$. Therefore $S_{high}$ is monotonically decreasing with $h$.
The complete proof is in Appendix A.1. Theorem 3.2 reveals the factors influencing the spectral distribution. Specifically, as the homophily level $h$ increases, the graph contains more low-frequency components. Moreover, the spectral distribution is related to both the number of edges $|\mathcal{E}|$ and the node distribution $\mathbb{E}(d_{inter})$, $\mathbb{E}(d_{intra})$.
Thus, graphs with different node distributions and homophily show spectral distribution diversity. We verify this using synthetic and real-world datasets. Using the Contextual Stochastic Block Model (CSBM) to generate graphs with homophily levels $0,0.2,\cdots,1.0]$, we observe their spectral distributions (details in Appendix B.4). We then calculate $S_{high}$ of all dimensions across CSBM graphs and show the Violin Plots in Figure 1(a). The figure shows that spectral distribution diversity exists among graphs, and validates Theorem 3.2 that higher homophily implies more low-frequency components. Also, spectral diversity exists across feature dimensions within the same graph due to the different feature distributions of these dimensions.
We also draw observations in real-world datasets with varying levels of homophily (dataset details in Appendix B.3 and Table 4) in Figure 1(b). Among them, Cora (McCallum et al. 2000), Citeseer (Sen et al. 2008) and Pubmed (Kipf and Welling 2016) are considered homophilic graphs, while Texas, Cornell, Wisconsin (Pei et al. 2020), Chameleon, and Squirrel (Rozemberczki, Allen, and Sarkar 2021) are heterophilic graphs. These observations align with our prior conclusions, confirming spectral distribution diversity across graphs and feature dimensions.

### v3

# Aligning the Spectrum: Hybrid Graph Pre-training and Prompt Tuning across Homophily and Heterophily

@@ Abstract
Graph “pre-training and prompt-tuning” aligns downstream tasks with pre-trained objectives to enable efficient knowledge transfer under limited supervision. However, current methods typically rely on single-filter backbones (e.g., low-pass), whereas real-world graphs exhibit inherent spectral diversity. Our theoretical Spectral Specificity principle reveals that effective knowledge transfer requires alignment between pre-trained spectral filters and the intrinsic spectrum of downstream graphs. This identifies two fundamental limitations: (1) Knowledge Bottleneck: single-filter models suffer from irreversible information loss by suppressing signals from other frequency bands (e.g., high-frequency); (2) Utilization Bottleneck: spectral mismatches between pre-trained filters and downstream spectra lead to significant underutilization of pre-trained knowledge. To bridge this gap, we propose HS-GPPT. We utilize a hybrid spectral backbone to construct an abundant knowledge basis. Crucially, we introduce Spectral-Aligned Prompt Tuning to actively align the downstream graph’s spectrum with diverse pre-trained filters, facilitating comprehensive knowledge utilization across both homophily and heterophily. Extensive experiments validate the effectiveness under both transductive and inductive learning settings.
@@ 1 Introduction
Graph “pre‑training and prompt tuning” Sun et al. (2022); Liu et al. (2023); Fang et al. (2024); Sun et al. (2023); Yu et al. (2024) has recently emerged as a powerful paradigm for handling data scarcity in graph‑based learning. By freezing a pre‑trained Graph Neural Network (GNN) and injecting small, learnable prompt structures (e.g., virtual nodes, subgraphs, or modified edges) into the input, one can adapt the model pretrained on graphs using self-supervision to new tasks on a target graph without updating its full parameter set. Under limited supervision, these prompts effectively bridge the gap between downstream objectives and the original pre‑training task, boosting knowledge transfer. Despite initial success, existing prompt tuning methods rely on the homophily assumption, learning node similarities in pre-training to empower downstream tasks. However, real-world graphs often contain heterophilic edges, meaning connections between dissimilar nodes, which leads to low homophily. Recent work Yu et al. (2024) shows that homophily-based prompt tuning methods struggle to generalize across graphs with varying homophily levels, under both transductive and inductive settings.
@@
To address this issue, we leverage the connection between homophily and spectral characteristics Chen et al. (2024a); Wan et al. (2024); Duan et al. (2024), and examine limitations of existing graph prompt tuning methods from a spectral perspective. As shown in Figure 1, we observe significant spectral distribution diversity across graphs with different homophily levels. Notably, lower homophily correlates with an increase in the high-frequency area $S_{high}$. However, current methods are typically restricted to a single kind of spectral knowledge (i.e., low-frequency information) with low-pass filters (e.g., GCN Kipf and Welling (2016)). Given this observation, we theoretically prove the Spectral Specificity principle: effective knowledge transfer requires alignment between pre-trained spectral filters and the intrinsic spectrum of downstream graphs. This reveals a twofold failure in existing methods: (1) Knowledge Bottleneck in the pre-training stage: single-filter models (e.g., low-pass filters) physically suppress high-frequency signals fundamental to heterophilic tasks, causing irreversible information loss; (2) Utilization Bottleneck in the prompt-tuning stage: even with a hybrid backbone, the inherent spectral diversity of downstream graphs often leads to partial alignment. Since each pre-trained filter captures a distinct frequency band, a specific downstream graph typically aligns with only a limited subset of filters, leaving the rich knowledge encoded in the remaining non-aligned filters underutilized.
@@
To address the heterophily issue in prompt tuning, we propose the Hybrid Spectral Graph Pre-training and Prompt Tuning model (HS-GPPT), which overcomes two challenges. The first challenge is: What spectral knowledge should be learned, and how? Our analysis shows that acquiring abundant spectral knowledge is essential for generalization. Therefore, we adopt a hybrid GNN backbone, where each filter captures a distinct spectral band. Building on prior methods Chen et al. (2024a) that handle only low- and high-pass filters, we extend the pre-training process to cover diverse frequency bands, constructing an abundant spectral basis.
@@
The second challenge is: How can prompts adaptively align with this abundant spectral knowledge to ensure full utilization? We introduce spectral-aligned prompt graphs: lightweight, learnable subgraphs that actively reshape the spectral distribution of downstream graphs to match the spectrum of the pre-trained filters. We theoretically prove the Spectral Adaptability of our prompts: for each frozen filter, a dedicated prompt can be optimized to transform the input spectrum. This strategy ensures that the abundant pre-trained spectral knowledge is fully utilized, maximizing the power of the frozen backbone.
@@
In summary, our contributions are: (i) The first theoretical analysis of hybrid spectral graph prompt tuning, proving the need for rich spectral pre-training and establishing spectral alignment principles; (ii) HS-GPPT, a novel framework leveraging abundant spectral knowledge for spectral downstream alignment, boosting knowledge transfer; (iii) Experiments across datasets with varying homophily show HS-GPPT outperforms baselines in both transductive and inductive settings, validating its generalization.
@@ 2 Related Work
Graph Prompt Tuning. Graph prompt tuning Sun et al. (2022); Fang et al. (2024); Sun et al. (2023); Liu et al. (2023); Yu et al. (2024) tackles data scarcity by leveraging frozen pre-trained knowledge through prompts. For example, GPrompt Liu et al. (2023) uses prompt vectors to unify pre-training and downstream tasks under a common template, while GPF Fang et al. (2024) inserts prompt nodes adaptable to various pre-training strategies. These methods, tailored to homophilic graphs, rely mainly on low-frequency signals and can be seen as special cases of our approach when restricted to low-frequency information. Our spectral analysis shows that under sparse supervision, large spectral gaps impede effective parameter learning, leading to the failure of existing homophily-based methods. Although ProNoG Yu et al. (2024), HeterGP Yan et al. (2025), and DAGPrompT Chen et al. (2025) have recently extended prompt tuning to heterophilic graphs via distribution-aware strategies, they operate in the spatial domain and rely on heuristic adaptations. Crucially, they overlook the fundamental spectral mismatch between the frozen backbone and diverse downstream graphs. In contrast, our method treats prompt tuning as a spectral alignment process, actively aligning the input spectrum to match pre-trained filters for better performance.
Heterophilic Graph Learning. Recent methods Xiao et al. (2024); Bo et al. (2021); Chen et al. (2024a); Wan et al. (2024) design GNN architectures and pre-training objectives for heterophilic graphs, such as leveraging high-frequency information Bo et al. (2021); Chen et al. (2024a); Wan et al. (2024) and discovering potential neighbors Jin et al. (2021); Pei et al. (2020). However, they rely on full-model fine-tuning. Without a mechanism to align pre-trained knowledge with downstream tasks, directly using their heterophilic priors can cause knowledge underutilization and even negative transfer, especially under limited supervision. Thus, these models and pre-training schemes for heterophilic graphs are ill-suited for prompt tuning methods.
@@ 3.1 Preliminaries
Recognizing that node classification is the main challenge on heterophilic graphs, we follow prior work Bo et al. (2021); Chen et al. (2024a) and select this task as our downstream objective. Specifically, the graph prompt tuning occurs under limited supervision, where for each class, only $K$ nodes are provided, also known as $K$-shot learning.
Homophily. Homophily refers to the tendency of nodes to connect with other nodes that are similar to themselves. Edge homophily level $h$ Zhu et al. (2020) measures the fraction of homophilic edges that connect nodes that share the same label, i.e., $h=\frac{|\{e_{uv}\in{\mathcal{E}}:y_{u}=y_{v}\}|}{|\mathcal{E}|}\in[0,1]$, where $|\mathcal{E}|$ denotes the number of edges and $y_{i}$ is the label of node $v_{i}$. A larger $h$ means a larger degree of homophily.
Spectral Graph Filters. Given a graph $\mathcal{G}={\{\mathcal{V},\mathbf{X},{\mathcal{E}}}\}$, let $\mathbf{A}\in\mathbb{R}^{N\times N}$ be the adjacency matrix and $\mathbf{D}\in\mathbb{R}^{N\times N}$ denote the diagonal degree matrix with $\mathbf{D}_{ii}=\sum_{j}\mathbf{A}_{ij}$. The graph Laplacian matrix $\mathbf{L}$ is $\mathbf{D}-\mathbf{A}$ (regular) or $\mathbf{I}-\mathbf{D}^{-\frac{1}{2}}\mathbf{A}\mathbf{D}^{-\frac{1}{2}}$ (normalized), where $\mathbf{I}$ is an identity matrix. $\mathbf{L}$ can be eigendecomposed as $\mathbf{L}=\mathbf{U}\mathbf{\Lambda}\mathbf{U}^{T}$, where $\mathbf{U}=[\mathbf{u}_{1},\mathbf{u}_{2},...,\mathbf{u}_{N}]$ are orthonormal eigenvectors and $\mathbf{\Lambda}=diag([\lambda_{1},\lambda_{2},...,\lambda_{N})]$ are corresponding eigenvalues. Typically, $\mathbf{U}$ is the graph Fourier basis and $\mathbf{\Lambda}$ are the frequencies. The objective of spectral graph filters is to design a function $g(\cdot)$ on $\mathbf{\Lambda}$ to learn the graph representation $\mathbf{Z}$:
@@
From a spatial view, different filtering characteristics aggregate diverse neighbor information. Low-pass filters capture node-neighbor similarity, while high-pass filters capture differences Luo et al. (2024).
Suppose the feature dimension $d=1$, the nodes’ feature is denoted as $\mathbf{x}\in\mathbb{R}^{N\times 1}$. The spectral graph signal $\mathbf{\hat{x}}$ is obtained by $\mathbf{\hat{x}}=(\hat{x}_{1},\hat{x}_{2},...,\hat{x}_{N})^{T}=\mathbf{U}^{T}\mathbf{x}$, where $\hat{x}_{i}=\mathbf{u}_{i}^{T}\mathbf{x}$ gives the projection of $\mathbf{x}$ to frequency $\lambda_{i}$. We denote the spectral energy as $\hat{x}_{k}^{2}/\sum_{i=1}^{N}\hat{x}_{i}^{2}$, which reflects the spectral distribution. The larger the $\hat{x}_{k}^{2}/\sum_{i=1}^{N}\hat{x}_{i}^{2}$ is, the more frequency components corresponding to $\lambda_{i}$ will be. To observe the spectral distribution intuitively, we introduce the high-frequency area  Tang et al. (2022).
@@
For a graph with spectral signal $\mathbf{\hat{x}}=(\hat{x}_{1},\hat{x}_{2},...,\hat{x}_{N})^{T}=\mathbf{U}^{T}\mathbf{x}$, the high-frequency area is calculated as $S_{high}=\frac{\sum_{k=1}^{N}\lambda_{k}\hat{x}_{k}^{2}}{\sum_{k=1}^{N}\hat{x}_{k}^{2}}=\frac{\mathbf{x}^{T}\mathbf{L}\mathbf{x}}{\mathbf{x}^{T}\mathbf{x}}$. $S_{high}$ quantifies the spectral distribution ($S_{high}\in[0,2]$). A larger $S_{high}$ indicates stronger high-frequency components, while a smaller implies greater low-frequency components.
@@ 3.2 Theoretical Analysis
In this section, we build the theoretical motivation for HS-GPPT. We first quantify the spectral distribution diversity, proving that structural shifts (e.g., homophily changes) inherently lead to significant spectral shifts. We then establish the spectral specificity principle: effective knowledge transfer requires strict alignment between pre-trained spectral filters and the downstream graph’s intrinsic spectrum. This highlights a dual necessity: constructing an abundant spectral basis (via pre-training) and employing spectral alignment (via prompt tuning) to ensure this knowledge is fully utilized.
3.2.1 Spectral Distribution Diversity
To understand the challenge of transferring knowledge across graphs, we first verify the connection between graph structure (homophily) and spectral properties. Given a graph $\mathcal{G}$ with edge set $\mathcal{E}$ and homophily level $h$, we define the expected normalized intra-class and inter-class distances as $\mathbb{E}(d_{intra})$ and $\mathbb{E}(d_{inter})$, respectively Ju et al. (2025) (details in Appendix A.1). Generally, connected nodes within the same class tend to have similar features, implying $\mathbb{E}(d_{inter})>\mathbb{E}(d_{intra})$. Based on the definition of high-frequency area $S_{high}$, we formalize the spectral-homophily correlation in Proposition 3.2.
Proposition 3.2 (Homophily-Spectrum Correlation).
The spectral energy distribution ($S_{high}$) is related to the homophily level $h$. Assuming bounded feature distances, $S_{high}$ exhibits a linear negative correlation with $h$:
$S_{high}=\mathcal{C}_{base}-\mathcal{C}_{gap}\cdot h,$ (3)
where $\mathcal{C}_{base}={|\mathcal{E}|}\mathbb{E}(d_{inter})$ and $\mathcal{C}_{gap}={|\mathcal{E}|}[\mathbb{E}(d_{inter})-\mathbb{E}(d_{intra})]$ are positive constants derived from the graph’s edge count and feature statistics.
The proof is provided in Appendix A.1. Proposition 3.2 reveals that spectral distributions are closely tied to homophily: a decrease in $h$ inherently leads to an increase in spectral energy $S_{high}$. This confirms the existence of spectral distribution diversity among downstream graphs with different homophily levels. We also validate this diversity using both synthetic Contextual Stochastic Block Models (CSBM) and real-world datasets. As shown in Figure 1(a) (details in Appendix B.4), as $h$ varies from 0 to 1, the spectral distributions (Violin Plots) exhibit clear diversity, perfectly aligning with the trend predicted in Eq. 3. Similarly, real-world datasets (Figure 1(b)) show distinct spectral patterns: homophilic graphs (e.g., Cora McCallum et al. (2000), Citeseer Sen et al. (2008)) are dominated by low-frequency components, while heterophilic ones (e.g., Texas Pei et al. (2020), Chameleon Rozemberczki et al. (2021)) show strong high-frequency energy.
This diversity poses a fundamental challenge: since standard pre-trained models are typically restricted to a single frequency band (mostly low-pass), transferring to diverse downstream tasks creates a significant spectral gap.